OpenStudy (anonymous):

What is the value of y in the parallelogram below? if necessary, round your answer to two decimal places

OpenStudy (anonymous):

@jim_thompson5910 @mathstudent55 @Mertsj @skullpatrol @RyanL.

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@marsss

OpenStudy (anonymous):

...........what grade are you in that you cant answer that?

OpenStudy (anonymous):

@Ahsher97 Wow you're nice -______- if you're not gonna help than don't come on here

zepdrix (zepdrix):



zepdrix (zepdrix):

Yah wow that's pretty rude... not everyone is learning on the same level as you ash :\

zepdrix (zepdrix):

Since it's a quadrilateral, how many degrees in total do the interior angles add up to? :)

OpenStudy (anonymous):

@zepdrix Awh, thanks! yeah, I'm just trying to finish up my math so I can graduate. It's all I have left, I just need a lil help! do they add up to 70 degrees? :)

zepdrix (zepdrix):

lol no silly! :) In a triangle, the interior angles add to \(180^o\). In a quadrilateral the interior angles add to \(360^o\). Since we have a PARALLELOGRAM, the angles opposite one another will be equal.

OpenStudy (anonymous):

i didnt post that sorry my friend is over :/

zepdrix (zepdrix):

If you look at the picture, that means that \(65^o=(2x+3)^o\) because those two angles are opposite one another. Ok ok but let's think a little bit differently, since we don't need \(x\). If that angle is 65, and the angle involving \(x\) is also 65, and we determined that a quadrilateral has 360 degrees total. Hmmmm. So those 2 angles \(65^o+65^o=130^o\) How much does that leave over for the other 2 angles? :D

OpenStudy (anonymous):

Oh! So what's the opposite? @Ahsher97 Lol, it's all good :)

zepdrix (zepdrix):

Oh thats weird :3 lol

OpenStudy (anonymous):

he shouldn't really talk he had to do summer school all four years of highschool

OpenStudy (anonymous):

Lolol awhh poor guy xD well at least you didn't say that! I bet your friend was just kiddin' around anyway :)

OpenStudy (anonymous):

@zepdrix is the opposite 65 degrees?

OpenStudy (anonymous):

since the left angle is opposite? :)

OpenStudy (tyteen4a03):

@Ahsher97 Please be nice to askers, regardless of which grade he/she is in. You have been there at one point. Another way to look at this is by using Interior Angles property (i.e 65+5y =180)

zepdrix (zepdrix):

This upper angle is also 65 since it's opposite the 65 in a parallelogram. How many degrees does that leave us with for the OTHER two angles if these two are each 65, and a quadrilateral has 360 degrees total? :D